OpenStudy (anonymous):

how do i solve -2 square root of (5x + 5) equals -30?

OpenStudy (anonymous):

first square both sides

OpenStudy (anonymous):

so is it (5x+5)^2 equals (-30)^2?

OpenStudy (anonymous):

\[-2\sqrt{5x+5}= -30\]

OpenStudy (anonymous):

what do i do with the -2?

OpenStudy (anonymous):

i'm lost

OpenStudy (anonymous):

I would divide both sides by -2 first so the numbers come out really nicely. Then square both sides.

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

\[-2\sqrt{5x+5} = -30\] \[\sqrt{5x+5} = 15\]

OpenStudy (anonymous):

Then square both sides.

OpenStudy (anonymous):

oh thank you :)